OpenStudy (anonymous):

center (-4,-11) and radius 7 units

OpenStudy (campbell_st):

use (x - h)^2 + (y -k)^2 = r^2 then expand and simplify

OpenStudy (anonymous):

(x+4)^2+(y+11)^2 = 49